From information on a previous question: The mean systolic blood pressure for a population of patients (µ) from a local clinic is 130 with a standard deviation (σ) of 18.

What is the probability that a patient chosen at random will have a systolic blood pressure between 93 and 126? Rounded to the nearest ten thousandth (4 places to the right of the decimal).
Answer #1

The following information has been provided:

u = 130, 0 = 18

We need to compute Pr(93 < X < 126 The corresponding z-values needed to be computed are:

2 = 1; -- _ 93 126 -20556

z) – 2 – + _ 126 – 130 = -0.222

Therefore, we get:

Pr(93 < X < 126) = Pr ( 937130 < z < 1207a 130 ) = Pr(-2.0556 < Z 5-0.2222) = Pr(Z < -0.2222) – Pr(Z < -2.0556) = 0.4121 – 0.

The following is obtained graphically:

normaldistributiongrapher.php?mean=130&s

Let me know in the comments if anything is not clear. I will reply ASAP! Please do upvote if satisfied!